OpenStudy (anonymous):

How do you do this ? Solve the equation: 4x2 = 36 A. {6, -6} B. {3, -3} C. {2/3, -2/3} D. {3/2, -3/2}

OpenStudy (callisto):

4x2 = 36 4x2 -36 =0 Factorise it, put each factor =0, and solve it, try ~

OpenStudy (anonymous):

B

OpenStudy (anonymous):

so 6,-6 ?

OpenStudy (callisto):

you also can take out the common factor first... i meant the '4'

OpenStudy (callisto):

nope..

OpenStudy (callisto):

factorise 4x2 -36 first

OpenStudy (anonymous):

so i need to find the GCF of 36 ?

OpenStudy (callisto):

36 and 4x^2

OpenStudy (anonymous):

4 .

OpenStudy (anonymous):

is the GCF

OpenStudy (callisto):

yes.. so you can rewrite it into 4(x^2-9), so now, factorise (x^2 -9)

OpenStudy (anonymous):

2/3 ?

OpenStudy (callisto):

nope...

OpenStudy (callisto):

factorise (x^2 -9)

OpenStudy (anonymous):

thats what i got. its either that or 3/2

OpenStudy (callisto):

\[x^2 -9 = (x-3)(x+3)\]

OpenStudy (anonymous):

so So if i do the square root of nine and times the answer two times that will give me my answer ?

OpenStudy (callisto):

4x2 -36 =0 4(x^2 -9) =0 4(x+3)(x-3) =0 (x+3)(x-3) =0 x+3 =0 or x-3 =0 x=-3 or x=3

OpenStudy (callisto):

no, \[\sqrt{9} \times 2 = 3\times2 = 6\]

OpenStudy (anonymous):

ok

OpenStudy (callisto):

do you understand?

OpenStudy (anonymous):

yes